OpenStudy (anonymous):

Which of the following expressions is equivalent to 6m - 10? 2(3m + 5) -2(5 - 3m) 2(3m - 10)

OpenStudy (anonymous):

@dtuniyants

OpenStudy (anonymous):

is it the last one?

Mehek (mehek14):

ok so 2 times what equals to 6 and -10

Mehek (mehek14):

no it is not the last one because 2*3 is 6 which is correct but 2* -10 is -20

OpenStudy (anonymous):

ohh so its the 1st one?

Mehek (mehek14):

try the first one 2*3m = 6m 2*5 = 10 it is supposed to be -10 not 10

Mehek (mehek14):

now the second one -2*5=-10 -2*-3m = 6 because when you multiply two negatives, the answe is a positive

Mehek (mehek14):

*6m*

Mehek (mehek14):

so what do you think now

OpenStudy (anonymous):

the second one

OpenStudy (shadowlegendx):

That's correct :)

OpenStudy (anonymous):

thnx

OpenStudy (shadowlegendx):

btw, you have an interesting smartscore ;P

Mehek (mehek14):

your welcome ^_^

OpenStudy (anonymous):

:D